Michael bought 2 notebooks and 1 pen for 7$. Rick bought 1 notebook and 5 pens for 8$. How much does the notebook cost ?
Answer by Cromlix(4381)

Hi there,
Make notebooks = x
Make pens = y
2x + y = 7.......(1)
x + 5y = 8.......(2)
Multiply Eq(1) by 5
10x + 5y = 35....(1)
x + 5y = 8.......(2)
Subtract Eq(2) from Eq(1)
9x = 27
x = 3
Substitute x = 3 into Eq(2)
x + 5y = 8
3 + 5y = 8
5y = 8 - 3
5y = 5
y = 1
Notebook costs $3
Pen costs $1
Hope this helps :-)